Given the relationship C = i_1/2/R ln (2), (A) wri
Answer #1

In problem given that,

t_{1/2}=0.0210\\ \delta t_{1/2}=\pm0.0005\\ R=98.5\\ \delta R=\pm0.1

Now C can be calculated directly as

C=\frac{t_{1/2}}{R\ln(2)}=\frac{0.0210}{98.5\times\ln(2))}=3.0758\times10^{-4}

To calculate \delta C, by taking the log on given equation and the by differentiation we get

\ln(C)=\ln(t_{1/2})-\ln(R)-\ln(\ln(2))

By differentiating, we get

\frac{\delta C}{C}=\frac{\delta t_{1/2}}{t_{1/2}}+\frac{\delta R}{R}+0\\ \Rightarrow \delta C=\left (\frac{0.0005}{0.0210}+\frac{0.1}{98.5} \right )\times C=\left (\frac{0.0005}{0.0210}+\frac{0.1}{98.5} \right )\times 3.0758\times10^{-4}=7.6356\times10^{-6}

therefore

C=3.0758\times10^{-4}\pm7.6356\times10^{-6}